# Federated attention contrastive learning models for prostate cancer diagnosis and Gleason grading

### v4

@@ 4.2 Experimental setups and evaluation metrics
We utilized the loss functions and algorithmic models outlined in Section 3.3 for each task. As an illustration, we conducted a simulated FL experiment employing 8 Tesla V100 GPUs, with each card representing an independent center. In this experiment, we employed a learning rate of 0.0002, an Adam optimizer, and a batch size of 1. To ensure optimal results, we trained the model for 60 epochs. Additionally, we implemented an early stopping strategy, halting training when the server model failed to demonstrate improvement after 20 consecutive epochs while ensuring a minimum of 40 epochs were completed. To promote convergence toward optimal outcomes, we calculated validation metrics for each epoch and selected the model with the highest performance on the validation set for the final evaluation of the test set. Our experiments encompassed three distinct settings: (1) Local Learning: Each center independently trained its data. (2) Global Learning: Aggregating datasets from all centers, we trained a multicenter centralized model. (3) Federated Learning: Training on data from all centers without sharing datasets, focusing solely on exchanging model parameters. By employing these strategies, we aimed to refine our approach and ensure a comprehensive evaluation of our model across different learning scenarios.
Table 4: The proportion of positive data across four centers.
@@
Table 5: The performance of the diagnostic model on the validation set was reported as the mean of five-fold cross-validation results. ($\mu$=0.1, $z$=0.1)
@@
FedAvg($z$) 0.9633 0.9248 0.9089 0.8855
FACL ($\mu$) 0.9626 0.9260 0.9096 0.8936
FACL ($\mu$, $z$) 0.9635 0.9256 0.9099 0.8859
@@
FedAvg($z$) 0.9708 0.9304 0.9160 0.8876
FACL ($\mu$) 0.9701 0.9343 0.9201 0.8976
FACL ($\mu$, $z$) 0.9715 0.9345 0.9204 0.8970
@@
FedAvg($z$) 0.9722 0.9341 0.9198 0.8977
FACL ($\mu$) 0.9731 0.9420 0.9291 0.9099
FACL ($\mu$, $z$) 0.9743 0.9447 0.9318 0.9202
@@
FedAvg($z$) 0.9682 0.9357 0.9210 0.9087
FACL ($\mu$) 0.9716 0.9435 0.9303 0.9201
FACL ($\mu$, $z$) 0.9718 0.9412 0.9269 0.9238
The distribution of data splits for the diagnostic task is outlined in Table 4. To assess the impact of data imbalance on the model’s performance, we partitioned the dataset from four centers (DiagSet-B-1, DiagSet-B-2, PANDA-1, PANDA-2), which contains a substantial number of samples, into positive proportions denoted as $\alpha$ (specifically, $\alpha={0.05,0.1,0.3,0.5}$). It is worth noting that we focused on conducting this verification solely within the realm of binary classification, primarily due to significant variations in the quantity and category of datasets. As such, we did not undertake a similar distribution experiment within the context of multi-category classification. Consequently, the simulation experiment of FL involved seven centers. Each center’s dataset was randomly divided into training and validation sets with an 80% and 20% ratio. Stratified sampling was employed to ensure a consistent class balance between each center’s training and validation sets. Furthermore, to evaluate the model’s robustness and generalization on previously unseen datasets, we utilized datasets from two independent centers for testing. In the case of the Gleason grading task, we implemented a FL model with six centers. We employed an independent dataset obtained from a private institution for testing purposes. On diagnostic and Gleason grading tasks, we evaluated the proposed FACL model using a variety of classification metrics, including AUC, F1 score, ACC, Recall, and Kappa score.
@@ 4.3 Results of diagnosis (Binary classification)
In the diagnostic task, we assessed models trained on datasets with varying degrees of heterogeneity, as indicated by the proportion factor $\bm{\alpha}=\left\{0.05,0.1,0.3,0.5\right\}$ in Table 4. Our experiments aimed to validate the performance of each local center model, FedAvg model, FACL model, and Centralized model across seven internal datasets and two external datasets (DiagSet-A and QHD).
@@ 4.3.1 Internal datasets
Table 6: Comparison results between FedAvg and FACL on DiagSet-A and QHD sets. ($\mu$=0.1, $z$=0.1)
@@
FedAvg ($z$) 0.9632 0.9156 0.9063 0.9559 0.9804 0.9491 0.9411 0.9699
FACL ($\mu$) 0.9666 0.9188 0.9110 0.9471 0.9825 0.9461 0.9372 0.9722
FACL ($\mu$,$z$) 0.9674 0.9148 0.9063 0.9471 0.9824 0.9379 0.9267 0.9769
@@
FedAvg ($z$) 0.9651 0.9135 0.9063 0.9251 0.9782 0.9553 0.9490 0.9630
FACL ($\mu$) 0.9678 0.9091 0.9016 0.9251 0.9758 0.9486 0.9411 0.9607
FACL ($\mu$,$z$) 0.9687 0.9164 0.9086 0.9427 0.9766 0.9466 0.9385 0.9630
@@
FedAvg ($z$) 0.9664 0.9121 0.9016 0.9603 0.9789 0.9490 0.9411 0.9676
FACL ($\mu$) 0.9710 0.9048 0.8946 0.9427 0.9801 0.9500 0.9424 0.9653
FACL ($\mu$,$z$) 0.9741 0.9063 0.8969 0.9383 0.9792 0.9239 0.9098 0.9676
@@
FedAvg ($z$) 0.9661 0.9012 0.8875 0.9647 0.9783 0.9478 0.9398 0.9653
FACL ($\mu$) 0.9718 0.8934 0.8782 0.9603 0.9796 0.9457 0.9372 0.9653
FACL ($\mu$,$z$) 0.9725 0.8915 0.8735 0.9779 0.9762 0.9170 0.9006 0.9699
The experimental results for the diagnosis task on the validation set are presented in Table 5, showcasing metrics such as AUC, F1, ACC, and Recall. Due to the different proportions of categories in the diagnostic task, the local center model’s overall performance is getting better with the increase of $\alpha$. When $\alpha$ is 0.05, the average ACC of 7 centers (Hebei-1, Hebei-2, Nanchang, DiagSet-B-1, DiagSet-B-2, PANDA-1, PANDA-2) model is only 0.8169, and the average AUC is only 0.9393. When $\alpha$ is 0.5, the average ACC of 7 centers (Hebei-1, Hebei-2, Nanchang, DiagSet-B-1, DiagSet-B-2, PANDA-1, PANDA-2) model reaches 0.8592, and the average AUC reaches 0.9499.
We observed superior performance in FL models compared to local center models. Specifically, the FACL model demonstrated a notable enhancement in AUC ranging from 0.02 to 0.04 when compared to individual centers. The deviation in results between the FACL model and the Centralized model is minimal, with less than a 0.01 difference in AUC. This outcome suggests that the FACL model has achieved an optimal value in line with federated learning theory, closely aligning with the performance of the Centralized model.
@@ 4.3.2 External datasets (DiagSet-A and QHD)
WSIs can vary widely in morphological appearance due to differences in institutional standards and protocols for tissue processing, slide preparation, and digitization of pathological images. Therefore, it is important to generalize the model application to actual clinical data. We validate the robustness of the model using two external datasets. We found that models trained using FL on multiple center data significantly improved the generalization performance of the models compared to models trained on single-center data.
The FedAvg and FACL model on the test set is presented in Table 6. The FACL model demonstrates superior AUC performance compared to the FedAvg model on both DiagSet-A and QHD datasets. On the DiagSet-A dataset, FACL ($\mu$,$z$) achieves optimal results, demonstrating its effectiveness. Additionally, on the QHD dataset, FACL ($\mu$) achieves optimal results without added noise. However, even when noise is introduced in FACL ($\mu$,$z$), the results remain highly satisfactory. The comparison results between the single-center model and the Centralized model can be found in Tables S1 and S2 in the appendix.
Table 7: The performance of the Gleason grading model on the validation set was reported as the mean of a five-fold cross-validation. ($\mu$=0.1, $z$=0.1)
@@
FedAvg ($z$) 0.8444 0.9136 0.6877 0.6951
FACL ($\mu$) 0.8535 0.9214 0.7201 0.7251
FACL ($\mu$, $z$) 0.8463 0.9126 0.6892 0.6969
4.4 Results of Gleason grading (Multiple classifications)
Given the scarcity of ISUP-annotated datasets, we assess the performance of local center models, the FedAvg model, the FACL model, and the Centralized model on five internal datasets and an external dataset (Nanchang).
4.4.1 Internal datasets
The validation set results for the Gleason scoring task are presented in Table 7. In comparison with models trained on single-center data, the FACL model exhibits significant improvements in Kappa score and AUC. The average Kappa across six centers (Hebei-1, Hebe-2, PANDA-1-1, PANDA-1-2, PANDA-2-1, PANDA-2-2) was 0.7379, whereas FACL achieved a Kappa of 0.8463. This highlights the effectiveness of federated learning in diagnosing prostate cancer across multiple categories. Notably, our proposed FACL model consistently outperforms the FedAvg model in Kappa score, regardless of the addition of noise ($z$). The Kappa scores for FACL ($\mu$) surpass FedAvg, and FACL ($\mu$,$z$) outperform FedAvg ($z$). This underscores the efficacy of our attention contrastive learning method.
4.4.2 External datasets (Nanchang)
Table 8: The Gleason grading model’s performance on Nanchang was reported as the five-fold mean.
@@ 4.4.2 External datasets (Nanchang)
The Kappa score exhibited a notable enhancement in the Gleason scoring task, as detailed in Table 8. Across the six centers, the average Kappa score reached 0.6437, while FedAvg achieved a Kappa score of 0.7247. Notably, FedAvg’s Kappa outperforms even the Centralized model, underscoring the robustness of models trained through federated learning. The comparative results of FedAvg and FACL models on the test set (Nanchang) are presented in Table 9. With the integration of an attention comparison mechanism, the FACL model demonstrates superior performance compared to FedAvg. FACL ($\mu$) achieves a Kappa score of 0.7298, surpassing FedAvg’s 0.7247, and FACL ($\mu$,$z$) at 0.7342 outperforms FedAvg ($z$) at 0.7227. This emphasizes the effectiveness of the attention contrastive learning mechanism introduced in the FACL model.
Table 9: Comparison results between FedAvg and FACL on the Nanchang set. ($\mu$=0.1, $z$=0.1)
@@
FedAvg($z$) 0.7227 0.8513 0.4448 0.4944
FACL ($\mu$) 0.7298 0.8451 0.4403 0.4883
FACL ($\mu$, $z$) 0.7342 0.8521 0.4376 0.4842
Fig. 4: Interpretability and visualization of FACL. Displayed are randomized selections of WSIs from external datasets, DiagSet-A and QHD, arranged in the first and second rows, respectively. The sequence progresses from left to right, showcasing the complete WSI, followed by its heatmap, a close-up of a local patch, and finally, the heatmap of the patch. The zoomed-in view of the local patch image underscores our model’s precise identification and representation of cancerous regions.
@@ 4.5 Interpretability and whole slide attention visualization
To assess the effectiveness of our proposed model in capturing morphological features of prostate cancer, a random subset of WSIs was selected from the test set. Heatmaps, as shown in Fig. 4, were generated using a patch size of $224\times 224$ and a 90% overlay. Each patch’s attention score was normalized to the range [0, 1], with high scores indicating regions crucial for diagnosis. To enhance clarity, RGB color maps were applied, with red representing high attention and blue indicating low attention. These heatmaps were then overlaid onto the original WSIs.
@@ 5 Conclusion
We proposed a novel FL model, referred to as FACL, designed for the diagnosis and Gleason grading of prostate cancer based on pathology images. We incorporated model attention comparison to enhance consistency between the server and client models, resulting in improved robustness and accuracy of the trained models. Additionally, to prevent malicious centers from inferring other centers’ data during the model training process and to safeguard the privacy of both model builders and data providers, we introduced noise to the model using differential privacy. This approach strikes a balance between privacy protection and model performance. While FACL has made strides in addressing privacy concerns for large-sized medical data, our work still has some limitations. First, we validated the model’s effectiveness only on prostate pathological images. The model’s performance can be further verified by testing on WSI datasets containing more diverse cancer types. Second, the number of private medical dataset centers we used remains relatively small. In the future, we aim to extend the application of FL to a wider range of medical diagnostic tasks in various medical scenarios. For instance, we envision its utilization in collaboration with medical institutions spanning different regions, facilitating FL in pathology image tasks and enabling a globally intelligent approach to rare disease diagnosis. Furthermore, as our current work has been conducted in a simulated environment, we recognize the need to consider communication efficiency in a real-world system. In a FL environment, training data remains distributed among numerous centers, each with potentially unreliable and relatively slow network connections. Overcoming the challenges of privacy, security, and communication efficiency associated with FL is crucial to drive further advancements in the healthcare industry.

### v5

@@ 4.2 Experimental setups and evaluation metrics
We utilized the loss functions and algorithmic models outlined in Section 3.3 for each task. As an illustration, we conducted a simulated FL experiment employing 8 Tesla V100 GPUs, with each card representing an independent center. In this experiment, we employed a learning rate of 0.0002, an Adam optimizer, and a batch size of 1. To ensure optimal results, the model was trained for over 60 epochs. Additionally, we implemented an early stopping strategy, halting training when the server model failed to demonstrate improvement after 20 consecutive epochs while ensuring that a minimum of 40 epochs were completed. To promote convergence toward optimal outcomes, we calculated the validation metrics for each epoch and selected the model with the highest performance in the validation set for the final evaluation of the test set.
Our experiments encompassed three distinct settings: (1) local learning: Each center independently trained its data; (2) global learning: aggregating datasets from all centers, along with training a multicenter centralized model; (3) federated learning: Training on data from all centers without sharing datasets, focusing solely on exchanging model parameters. To comprehensively validate the effectiveness of attention-consistency learning, we introduced a noisy model, FedAvg-N, as another baseline. Noise introduces significant interference into the model, and we aim to ascertain whether attention consistency still holds under noisy conditions. By employing these strategies, we aim to ensure a comprehensive evaluation of our model across different learning scenarios.
Table 4: The proportion of positive data across four centers
@@
The distribution of data splits for the diagnostic task is presented in Table 4. To assess the impact of data imbalance on the model’s performance, we partitioned the dataset from four centers (DiagSet-B-1, DiagSet-B-2, PANDA-1, PANDA-2), which contains a substantial number of samples, into positive proportions denoted as $\alpha$ (specifically, $\alpha\in\{0.05,0.1,0.3,0.5\}$). Notably, we focused on conducting this verification solely within the realm of binary classification, primarily because of significant variations in the quantity and category of the datasets. Therefore, we did not conduct a similar distribution experiment in the context of multiclass classification. Consequently, the FL simulation experiment involved seven centers. Each center’s dataset was randomly divided into training and validation sets at a ratio of 80% and 20%, respectively. Stratified sampling was employed to ensure a consistent class balance between each center’s training and validation sets. Furthermore, we used datasets from two independent centers to evaluate the robustness and generalization of the model to previously unseen datasets for testing. For the Gleason grading task, we implemented an FL model with six centers. Additionally, we employed an independent dataset obtained from a private institution for testing. On diagnostic and Gleason grading tasks, we evaluated the proposed FACL model using a variety of classification metrics, including area under the curve (AUC), F1 score, ACC, Recall, and Kappa score.
Table 5: The performance of the diagnostic model on the validation set was reported as the mean of five-fold cross-validation results
@@
FedAvg-$N$ 0.9633 0.9248 0.9089 0.8855
FACL 0.9626 0.9260 0.9096 0.8936
FACL-$N$ 0.9635 0.9256 0.9099 0.8859
@@
FedAvg-$N$ 0.9708 0.9304 0.9160 0.8876
FACL 0.9701 0.9343 0.9201 0.8976
FACL-$N$ 0.9715 0.9345 0.9204 0.8970
@@
FedAvg-$N$ 0.9722 0.9341 0.9198 0.8977
FACL 0.9731 0.9420 0.9291 0.9099
FACL-$N$ 0.9743 0.9447 0.9318 0.9202
@@
FedAvg-$N$ 0.9682 0.9357 0.9210 0.9087
FACL 0.9716 0.9435 0.9303 0.9201
FACL-$N$ 0.9718 0.9412 0.9269 0.9238
@@ 4.3 Results of diagnosis (Binary classification)
In the diagnostic task, we assessed models trained on datasets with varying degrees of heterogeneity, as indicated by the proportion factor $\bm{\alpha}\in\left\{0.05,0.1,0.3,0.5\right\}$ in Table 4. Our experiments aimed to validate the performance of each local center, FedAvg, FACL, and Centralized model across seven internal datasets and two external datasets (DiagSet-A and QHD).
@@ 4.3.1 Internal datasets
Table 6: Comparison results between FedAvg and FACL on DiagSet-A and QHD sets
@@
FedAvg-$N$ 0.9632 0.9156 0.9063 0.9559 0.9804 0.9491 0.9411 0.9699
FACL 0.9666 0.9188 0.9110 0.9471 0.9825 0.9461 0.9372 0.9722
FACL-$N$ 0.9674 0.9148 0.9063 0.9471 0.9824 0.9379 0.9267 0.9769
@@
FedAvg-$N$ 0.9651 0.9135 0.9063 0.9251 0.9782 0.9553 0.9490 0.9630
FACL 0.9678 0.9091 0.9016 0.9251 0.9758 0.9486 0.9411 0.9607
FACL-$N$ 0.9687 0.9164 0.9086 0.9427 0.9766 0.9466 0.9385 0.9630
@@
FedAvg-$N$ 0.9664 0.9121 0.9016 0.9603 0.9789 0.9490 0.9411 0.9676
FACL 0.9710 0.9048 0.8946 0.9427 0.9801 0.9500 0.9424 0.9653
FACL-$N$ 0.9741 0.9063 0.8969 0.9383 0.9792 0.9239 0.9098 0.9676
@@
FedAvg-$N$ 0.9661 0.9012 0.8875 0.9647 0.9783 0.9478 0.9398 0.9653
FACL 0.9718 0.8934 0.8782 0.9603 0.9796 0.9457 0.9372 0.9653
FACL-$N$ 0.9725 0.8915 0.8735 0.9779 0.9762 0.9170 0.9006 0.9699
The experimental results for the diagnosis task on the validation set are presented in Table 5, demonstrating metrics such as AUC, F1, ACC, and Recall. As $\alpha$ increases, the overall performance of the local center model improves due to the different proportions of categories in the diagnostic task. When $\alpha$ is 0.05, the average ACC of the models from seven centers (Hebei-1, Hebei-2, Nanchang, DiagSet-B-1, DiagSet-B-2, PANDA-1, PANDA-2) is only 0.8169, with an average AUC of 0.9393. However, when $\alpha$ is 0.5, the average ACC of these models (Hebei-1, Hebei-2, Nanchang, DiagSet-B-1, DiagSet-B-2, PANDA-1, PANDA-2) reaches 0.8592, accompanied by an average AUC of 0.9499.
We observed superior performance in the FL models compared to the local center models. Specifically, the FACL model demonstrated a notable enhancement in AUC, ranging from 0.02 to 0.04 when compared to individual centers. The deviation in the results between the FACL model and the Centralized model is minimal, with less than a 0.01 difference in AUC. This outcome suggests that the FACL model achieved an optimal value in line with federated learning theory, closely aligning with the performance of the Centralized model.
@@ 4.3.2 External datasets (DiagSet-A and QHD)
WSIs can vary widely in morphological appearance due to differences in institutional standards and protocols for tissue processing, slide preparation, and digitization of pathological images. Therefore, it is important to generalize the application of the model to actual clinical data. We validate the robustness of the model using two external datasets. Our findings indicate that models trained using FL on multicenter data significantly improved the generalization performance of the models compared to models trained on single-center data.
The FedAvg and FACL models for the test sets are listed in Table 6. The FACL model demonstrated superior AUC performance compared with the FedAvg model for both the DiagSet-A and QHD datasets. Similarly, the FACL-$N$ model outperformed the FedAvg-$N$ model in terms of AUC performance. For the DiagSet-A dataset, FedAvg-$N$ achieved optimal results, demonstrating its effectiveness. In addition, on the QHD dataset, the FACL achieved optimal results without added noise. However, the results remained highly satisfactory even when noise was introduced into the FACL-$N$ model. The comparison results between the single-center and Centralized models can be found in the tables S1 and S2 in the appendix.
Table 7: The performance of the Gleason grading model on the validation set was reported as the mean of a five-fold cross-validation
@@
FedAvg-$N$ 0.8444 0.9136 0.6877 0.6951
FACL 0.8535 0.9214 0.7201 0.7251
FACL-$N$ 0.8463 0.9126 0.6892 0.6969
Table 8: The Gleason grading model’s performance on Nanchang was reported as the five-fold mean
@@
4.4 Results of Gleason grading (Multiple classifications)
Given the scarcity of ISUP-annotated datasets, we assessed the performance of the local center models, FedAvg model, FACL model, and centralized model on five internal datasets and an external dataset (Nanchang).
4.4.1 Internal datasets
Fig. 4: Interpretability and visualization of FACL. Randomized selections of WSIs from external datasets, DiagSet-A and QHD, are arranged in the first and second rows, respectively. The sequence progresses from left to right, showcasing the complete WSI, followed by its heatmap, a close-up of a local patch, and finally the heatmap of the patch. The zoomed-in view of the local patch image indicates our model’s precise identification and representation of cancerous regions.
The validation results for the Gleason scoring task are listed in Table 7. Compared to models trained on single-center data, the FACL model exhibited significant improvements in the Kappa score and AUC. The average Kappa across the six centers (Hebei-1, Hebe-2, PANDA-1-1, PANDA-1-2, PANDA-2-1, and PANDA-2-2) was 0.7379, whereas FACL achieved a Kappa of 0.8463. This highlights the effectiveness of federated learning for prostate cancer diagnosis across multiple categories. Notably, the proposed FACL model consistently outperformed the FedAvg model in terms of the Kappa score, regardless of the addition of noise ($N$). The Kappa scores for the FACL surpassed those for the FedAvg, and the FACL-$N$ outperformed the FedAvg-$N$. This underscores the efficacy of the proposed attention-consistent learning method.
4.4.2 External datasets (Nanchang)
Table 9: Comparison results between FedAvg and FACL on the Nanchang set
@@ 4.4.2 External datasets (Nanchang)
FedAvg-$N$ 0.7227 0.8513 0.4448 0.4944
FACL 0.7298 0.8451 0.4403 0.4883
FACL-$N$ 0.7342 0.8521 0.4376 0.4842
The Kappa score exhibited a notable enhancement in the Gleason score, as shown in Table 8. Across the six centers, the average Kappa score reached 0.6437, whereas the FedAvg achieved a Kappa score of 0.7247. Notably, FedAvg’s Kappa outperformed that of the Centralized model, underscoring the robustness of models trained through federated learning.
The comparative results of the FedAvg and FACL models for the test set (Nanchang) are presented in Table 9. With the integration of an attention-consistent mechanism, the FACL model demonstrated superior performance compared with FedAvg. FACL achieved a Kappa score of 0.7298, surpassing FedAvg’s 0.7247, and FACL-$N$ at 0.7342 outperforms FedAvg-$N$ at 0.7227. This emphasizes the effectiveness of the attention-consistent learning mechanism introduced in the FACL model.
@@ 4.5 Interpretability and whole slide attention visualization
To assess the effectiveness of the proposed model in capturing the morphological features of prostate cancer, a random subset of the WSIs was selected from the test set. Heatmaps, as shown in Fig. 4, were generated using a patch size of $224\times 224$ and 90% overlay. The attention score of each patch was normalized to the range [0, 1], with high scores indicating regions crucial for diagnosis. RGB color maps were used to enhance the clarity, with red representing high attention and blue indicating low attention. These heat maps were overlaid onto the original WSIs.
@@ 5 Conclusion
We propose a novel FL model, referred to as the FACL, designed for diagnosis and Gleason grading of prostate cancer based on pathological images. We incorporated attention consistency to enhance consistency between the server and client models, resulting in improved robustness and accuracy of the trained models. Additionally, to prevent malicious centers from inferring the data of other centers during the model-training process and to safeguard the privacy of both model builders and data providers, we introduced noise into the model using differential privacy. This approach strikes a balance between privacy protection and performance. Although the FACL has made strides in addressing privacy concerns for large-sized medical data, our work still has some limitations. First, the effectiveness of the model was validated using only pathological prostate images. The performance of the model was further verified by testing WSI datasets containing more diverse cancer types. Second, the number of private medical dataset centers used remained relatively small. Future research aims to extend FL applications to various medical diagnostic tasks in diverse scenarios. Collaboration with medical institutions globally could enhance FL in pathology image tasks and enable intelligent approaches to rare disease diagnosis. Furthermore, acknowledging that our current work was conducted in a simulated environment, we recognize the need to consider communication efficiency in real-world systems. In an FL environment, training data remain distributed among numerous centers, each with potentially unreliable and relatively slow network connections. Overcoming the challenges of privacy, security, and communication efficiency associated with FL is crucial for driving further advancements in the healthcare industry.